OpenStudy (anonymous):

What is the solution of the system of equations? y= -3x+8 y= -5x - 2

OpenStudy (anonymous):

first solve one equation for x and then plug it into the other

OpenStudy (radar):

y=y -3x+8=-5x-2 2x = -2-8 2x = -10 x=-5 Now solve for y.

OpenStudy (anonymous):

what he said

OpenStudy (anonymous):

23?

OpenStudy (radar):

You did good.

OpenStudy (anonymous):

(-5,23)